OpenStudy (anonymous):

Which of the following would eliminate the variable on the left side of the given equation? 36 - 19w = -6w + 41

OpenStudy (anonymous):

If you want to eliminate the variable w on the left side of the equation, then you add 19w to both sides of the equation. So you have 36+(19w-19w)=(19w-6w)+41 this becomes 36+0=13w+41 which becomes 36=13w+41